the maganetic field produced by a soenoid in a MRI system desigend for measurements on whole human bodies has a field strength of 7T, and the current in the solenoid is 200 A. what is the number of turns per meter of length of solenoid?
Answer #1

magnetic field of solenoid, B = u*n*I (where u = permeability of the medium, n = no. of turns per meter, I = current)

n = B/(u*I)= 7/(200*4*pi*10^-7) = 27852

no. of turns per meter =  27852